OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -12, -7, -2, ..

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@ParthKohli

OpenStudy (trojanpoem):

-12 - (-17) = 5 nth trm = a0 + m(n-1) a0 is the first in the sequence m is the difference ( 5) n is the place

OpenStudy (anonymous):

a_n = -17 + 5 (n -1)